# Over-parameterization and Adversarial Robustness in Neural Networks: An Overview and Empirical Analysis

### v1

@@ Abstract
Thanks to their extensive capacity, over-parameterized neural networks exhibit superior predictive capabilities and generalization. However, having a large parameter space is considered one of the main suspects of the neural networks’ vulnerability to adversarial examples— input samples crafted ad-hoc to induce a desired misclassification. Relevant literature has claimed contradictory remarks in support of and against the robustness of over-parameterized networks. These contradictory findings might be due to the failure of the attack employed to evaluate the networks’ robustness. Previous research has demonstrated that depending on the considered model, the algorithm employed to generate adversarial examples may not function properly, leading to overestimating the model’s robustness. In this work, we empirically study the robustness of over-parameterized networks against adversarial examples. However, unlike the previous works, we also evaluate the considered attack’s reliability to support the results’ veracity. Our results show that over-parameterized networks are robust against adversarial attacks as opposed to their under-parameterized counterparts.
@@
keywords:
journal: Pattern Recognition Letters
@@ 1 Introduction
Neural networks have reported outstanding performances in the last decade. This advancement has been possible not only thanks to the availability of large datasets but also to the large number of parameters employed. The networks that have reported better performances are often over-parameterized [1], which means that their capacity is higher than the one needed to interpolate the training data (to fit them perfectly).
In modern Machine Learning (ML) practices, rich function classes like neural networks (NN) are highly capable of training models that generalizes well on test samples, meaning giving accurate predictions on samples not seen during the training, provided the test data is sampled from the same distribution as training data [2, 3]. However, this phenomenon is at odds with the classical understanding of bias-variance tradeoff that predicts lower bias in increasingly complex [^1] The classical bounds in statistical learning theory point to the size of the model in terms of the number of parameters as a measure of complexity. models at the expense of higher variance typically supporting parsimonious networks. Such a mismatch in theoretical and empirical understanding has led the community to study the behaviour of over-parameterized networks deeply. Belkin et al.’s theory of double-descent curve (Figure 1) confirms that the variance and, subsequently, the risk of the model decreases as the size of the model goes beyond the interpolation point. It is even considered that the lowest generalization error is expected at the infinitely large model.
Large neural networks have shown impressive performance at test time across domains, leading to their widescale applicability. However, surprisingly, these models are vulnerable to adversarial examples: input samples carefully perturbed to be classified as the attacker desires [4, 5, 6]. Computer vision poses a particularly challenging issue since it has been shown that for the attackers, it is sufficient to apply a tiny perturbation, imperceptible to human eyes, to have a sample misclassified with high confidence  [7, 8]. For security-sensitive applications such as autonomous cars and face recognition, it is imperative that the models are robust against adversarial examples while performing well on natural samples.
From a classical understanding of the optimal size of neural network, most neural networks, to minimize the empirical risk, were biased to choose the smallest possible model that fits the training data well, i.e., stay on the left side of the double descent curve. Hence, most adversarial evaluations are performed on these under-parameterized networks.
Recently, given the massive performance hike of over-parameterized models and their generalizable properties, their robustness to adversarial examples has been investigated in multiple works. However, the claim about the security of these models are contradictory, and the relationship of over-parameterized models with adversarial robustness is unclear. Recent empirical studies showed their near-optimal results even when the neural network and kernel machines trained to interpolation are corrupted with noise  [9, 10]. Adversarial examples, different from random noise, are carefully crafted samples designed to intentionally target the vulnerability of the model.
One of the major reasons for these contradictory results may be the unreliable evaluation of adversarial robustness. In various recent works where a robustness evaluation was performed, adversarial attacks were found to have failed [11], i.e., the attacks generated were ineffective and hence useless for evaluation, giving a false sense of security  [12]. However, the articles assessing the robustness of overparametrized networks with respect to their non-overparametrized counterpart have never investigated the effectiveness of the attack employed to perform the robustness evaluation.
We evaluate the adversarial robustness of different models, trained on a spectrum of model capacity against Projected Gradient Descent (PGD) attacks. To be consistent with the best practices and evaluate robustness reliably we test the models on diverse-attack framework like AutoAttack. To avoid attack failure, we also used tools to verify the attack effectiveless. The empirical results from the experiments suggest that highly parameterized models are more robust as compared to under-parameterized models.
We discuss the background and the notations in section 2, the methodology adopted to conduct experiments in section 3, and experimental setup and results in section 4. We conclude by discussing the major takeaways and future research directions in section 5.
Figure 1: The image is taken from the Belkin et al.  [1] where they reconcile the modern understanding of bias-variance with conventional wisdom: a) the classical U-shape risk curve derived from bias-variance trade-off, b) modern double-descent risk curve where risk depreciate beyond the point of interpolation threshold, a.k.a. in the over-parameterization regime
@@ 2.1 Over-parameterization and Double-descent curve
When designing a model architecture to solve an ML problem, from classical understanding, we choose a model that is rich enough to capture the underlying feature space without fitting too closely with the training data (or overfitting) and simple enough to avoid not being able to learn patterns in the data (or underfitting). The balance between underfitting and overfitting is widely known as bias-variance tradeoff plotted on a risk capacity curve (Figure 1(a)). From this understanding, we would want to find a ”sweet spot” where the model has enough capacity to perform well on test data while keeping the training error low. Having a large model with zero training error is conventionally avoided, as stated by Trevor Hastie in  [13], ”a model with zero training error is overfit to the training data and will typically generalize poorly” which is still a widely accepted notion.
However, large NN are highly expressive and can efficiently optimize many parameters to fit the training data. Large number of parameters of the neural network, traditionally considered a case of overfitting, are better suited for generalization  [14]. Further, Belkin et al.’s  [1] who coined the term double-descent curve (Figure 1(b)) suggested that when training a model on the increasing function capacity (referred to the size of network) until the point of zero training error/interpolation, typically have high risk and poor generalization error. However, training beyond the interpolation point (in over-parameterized regime), it is observed that the variance of the model decreases which leads to the overall decrease in the risk, giving improved generalization performance. Further work by Nakkiran et al. [15] showed that this generalization trend is not just a function of model capacity but also a function of number of epochs in training. In some literature, it is also argued that good generalization performance is due to the inductive bias of gradient-based algorithms, e.g.,  [16, 17].
@@ 2.2 Adversarial Robustness
[…]
@@ 2.3 Adversarial Robustness of Over-parameterized Network
While a considerable number of works [18, 19, 20, 21, 22] assess the robustness of adversarially trained networks, which is out of the scope of this article, the findings of various related works suggest that overparameterization would decrease robustness. In the seminal work by Fu et al. [23], the authors claim that decreasing the number of nodes reduces the sensitivity to small perturbations of the input samples. Sagawa et al. [24] suggested that the over-parameterization can cause the exacerbation of spurious correlation, i.e., the relation between the biased features of input like background, texture, or unimportant pixels with the labels. However, in [25], the authors showed that wider neural networks tend to have better robustness, assessed with the PGD attack. The results in this paper were also replicated in [26] and are supported by [27], where the authors claim that obtaining robustness may require exponentially more parameters than obtaining accuracy on the clean test data. The contradiction present in the literature has been highlighted in [28], which provides a theoretical analysis that only encompasses two-layer neural networks, and in [29], which, with an analysis limited to linear models, shows that overparametrization may be either a source of additional robustness or brittleness. The analysis performed on deep neural networks by the works affirming that overparameterization increases robustness [25, 26] is empirical and based on the PGD attack. Therefore, it might have failed [30, 31], providing unreliable results that have caused a false sense of robustness for the overparameterized networks. The goal of this work is to verify if this is the reason that led to the abovementioned contradiction between the state-of-the-art works.
Figure 2: Base performance of trained classifiers against the parameter size: (a) CNN trained on MNIST, (b) linear classifier trained on MNIST, and (c) ResNet model trained on CIFAR10
@@ 3.1 Notations
We consider a data space [$\mathcal{X}$, $\mathcal{Y}$] where $\mathcal{X}$ is a $d$-dimensional feature space of input samples and $\mathcal{Y}$ is a label space with c-classes such that $\mathcal{Y}\in[c]$. The training data is represented as $\mathcal{D}=(\mathbf{x}_{i},y_{i})_{i=1}^{n}$ where $n$ is the size of the training set. We use $L(\mathcal{D},\theta)$ to represent loss incurred by classifier $f\in\mathcal{H}:\mathcal{X}\rightarrow\mathcal{Y}$ where $\mathcal{H}$ is a class of function and $\theta$ are the model parameters. Typically the loss is computed by averaging the loss function $l(\mathbf{x},y,\theta)$ over the training set, i.e., $\frac{1}{n}\sum_{i=1}^{n}l(\mathbf{x}_{i},y_{i},\theta)$. We assume that the classifier $f$ is learned by minimizing the loss function and best function class $f^{*}$ is selected using empirical risk minimization:
@@
Figure 3: SEC plots of over-parameterized models using PGD $l_{2}$ attack on (a) CNN trained on MNIST, (b) FC-ReLU trained on MNIST, and (c) ResNet trained on CIFAR10
@@ 3.2 Attack Approach
The attack approaches considered in our experiments have the following threat model: The attacker’s main goal is to cause integrity violation to evade detection without compromising normal operation using indiscriminate attack whose goal is to have the test-instance misclassified as any class except the true one. The attacker’s capability is limited to influencing data samples at test time to evade the trained model, making it an evasion attack. We consider both white-box setting, where the attacker has complete access to both model and data distribution, and the black-box setting, where the attacker has no access to either data or distribution.
To evaluate the security of over-parameterized models, we use the widely-used white-box attack approach, which generates attacks on the known trained model parameters.
PGD is a widely used evasion attack algorithm to evaluate the robustness of a model and is a good starting point for security evaluation. After evaluating on PGD, as explained in later sections, we perform experiments on sophisticated diversified ensemble attack: AutoAttack.
3.2.1 Projected Gradient Descent: PGD
PGD is an iterative optimization-based attack [32]. It is widely used for the evaluation of the models’ robustness. Given a sample $\bm{x}$, the attacker uses the model parameters w and dataset $\mathcal{D}$ to generate the adversarial example in the feasible domain $\Gamma={\{\mathbf{x}_{t}\in\mathcal{X}:||\mathbf{x}_{t}-\mathbf{x}_{0}||_{p}<\epsilon\}}$:
$\mathbf{x}_{t+1}=\prod_{\Gamma}(\mathbf{x}_{t}+\alpha\,sgn(\nabla_{x}\,L(\mathcal{D},\theta))$
where $\nabla$ denotes the gradient operator, $\epsilon$ is the attack budget that defines the strength of the attack and $\alpha$ denote the size of the current step. This procedure is usually repeated many times to find the adversarial examples.
@@ 3.2.2 AutoAttack
AutoAttack is an ensemble of different typologies of attacks [33]. It leverages three white-box and one black-box attack. The three white-box attacks include Automated PGD with Carlini-Wagner loss $APGD_{CW}$, Automated PGD with Difference of Logits Ratio loss $APGD_{DLR}$, targeted Fast Adaptive Boundary (FAB) attack, and Square Attacks as the black-box attack  [34]. The Automated PGD (APGD) is the standard PGD attack with a fixed number of iterations and adaptive step size. It uses Carlini-Wagner Loss and Difference of Logits Ratio loss instead of cross-entropy loss.
$APGD_{CW}$ is an untargeted attack whereas $APGD_{DLR}$ is a targeted attack. Namely, it computes the adversarial example ad-hoc to have it misclassified as a specific class. If it does not succeed, it considers a different class as the target class. The third parameter-free white-box attack is the targeted version of the Fast Adaptive Boundary (FAB) attack  [35]. Finally, Square Attack is an untargeted black-box attack based on random search and is not affected by gradient masking. Gradient masking is a common adversarial defense against gradient-based attack algorithms that makes the gradient usually computed by these attacks uninformative, which makes the attacker unable to compute an effective adversarial example. However, thanks to the transferability property of adversarial examples [36], these defense techniques can be easily broken by computing an attack against a model owned by the attacker (surrogate). These four attacks are packaged in a single module that serves as a diversified attack and helps evaluate the reliability of the results we obtained from our primary PGD experiments.
@@ 3.3 Evaluation Procedure
To evaluate the security of neural networks, we take into account different neural network architectures. These architectures are then scaled with different numbers of hidden units in their layers as done in [37]. We consider a model architecture, expanded to ten models ranging from under-parameterized to over-parameterized regime $\theta\in\{\theta_{1},\theta_{2},..,\theta_{m}\}$. The models are first trained on the training data. Then, for each model $\theta_{i}$, its base performance is evaluated using benign test data, i.e., where there is no perturbation ${\epsilon_{0}}$. Consequently, we generate the adversarial examples ad-hoc against the considered model $\theta_{i}$ for varying degrees of attack budgets $\epsilon\in\{\epsilon_{1},\epsilon_{2}..,\epsilon_{k}\}$ and for each perturbation budget we compute the Error Rate (namely the percentage of misclassified samples) of $\theta_{i}$ on them. These error rates are then used to generate the abovementioned SEC curve that shows how the error rate of the model $\theta_{i}$ increases for a growing perturbation budget of the attacker.
Figure 4: SEC plots of over-parameterized models using AutoAttack on (a) MNIST trained on CNN, (b) MNIST trained on FC-ReLU, and (c) CIFAR10 trained on ResNet

### v2

@@ Abstract
Thanks to their extensive capacity, over-parameterized neural networks exhibit superior predictive capabilities and generalization. Nevertheless, having a large parameter space is considered one of the main suspects of the neural network vulnerability to adversarial examples— input samples crafted ad-hoc to induce a desired misclassification. Relevant literature has claimed contradictory remarks in support of and against the robustness of over-parameterized networks. These contradictory findings might be due to the failure of the attack employed to evaluate the robustness of the networks. Previous research has demonstrated that depending on the considered model, the algorithm employed to generate adversarial examples may not function properly, leading to overestimating its robustness. In this work, we empirically study the robustness of over-parameterized networks against adversarial examples. However, unlike the previous works, we also evaluate the considered attack’s reliability to support the results’ veracity. Our results show that over-parameterized networks are robust against adversarial attacks as opposed to their under-parameterized counterparts.
@@
 [^1] This manuscript is under review at the Discover Artificial Intelligence.
@@ 1 Introduction
Neural networks have reported outstanding performances in the last decade. This advancement has been possible not only thanks to the availability of large datasets but also to the large number of parameters employed. The networks that have reported better performances are often over-parameterized [1]: the number of learned parameters is larger than the number of training samples. In modern Machine Learning (ML) practices, rich function classes like neural networks (NN) are highly capable of training models that generalizes well on test samples, meaning giving accurate predictions on samples not seen during the training, provided the test data is sampled from the same distribution as training data [2, 3].
Figure 1: Risk curves. The bias is represented in blue and the variance is represented in red. Left plot: Double-descent risk curve; Right plot: The risk when, as for deep neural networks, the bias dominates.
However, this phenomenon is at odds with the classical understanding of bias-variance tradeoff that predicts lower bias in increasingly complex [^2] The classical bounds in statistical learning theory point to the size of the model in terms of the number of parameters as a measure of complexity. models at the expense of higher variance, typically supporting parsimonious networks. Such a mismatch in theoretical and empirical understanding has led the community to study the behavior of over-parameterized networks deeply. Belkin et al.’s theory of double-descent curve (Figure 1) confirms that the variance and, subsequently, the risk of the model decreases as the size of the model goes beyond the interpolation point. It is even considered that the lowest generalization error is expected at the infinitely large model. Large neural networks have shown impressive performance at test time across domains, leading to their widescale applicability. However, surprisingly, these models are vulnerable to adversarial examples: input samples carefully perturbed to be classified as the attacker desires [4, 5, 6]. Computer vision poses a particularly challenging issue since it has been shown that for the attackers, it is sufficient to apply a tiny perturbation, imperceptible to human eyes, to have a sample misclassified with high confidence  [7, 8]. For security-sensitive applications such as autonomous cars and face recognition, it is imperative that the models are robust against adversarial examples while performing well on natural samples. From a classical understanding of the optimal size of the neural network, most neural networks, to minimize the empirical risk, were biased to choose the smallest possible model that fits the training data well, i.e., stay on the left side of the double descent curve. Hence, most adversarial evaluations are performed on these under-parameterized networks. Recently, given the massive performance hike of over-parameterized models and their generalizable properties, their robustness to adversarial examples has been investigated in multiple works. However, the claims about the security of these models are contradictory, and the relationship of over-parameterized models with adversarial robustness is unclear. Recent empirical studies showed their near-optimal results even when the neural network and kernel machines trained to interpolation are corrupted with noise  [9, 10]. Adversarial examples, distinct from random noise, are carefully crafted samples designed to intentionally target the vulnerability of the model. One of the major reasons for these contradictory results may be the unreliable evaluation of adversarial robustness. In various recent works where a robustness evaluation was performed, adversarial attacks were found to have failed [11], i.e., the attacks generated were ineffective and hence useless for evaluation, giving a false sense of security  [12]. However, the articles assessing the robustness of overparameterized networks with respect to their non-overparameterized counterpart have never investigated the effectiveness of the attack employed to perform the robustness evaluation. We evaluate the adversarial robustness of different models, trained on a spectrum of model capacity against Projected Gradient Descent (PGD) attacks. To be consistent with the best practices and evaluate robustness reliably, we test the models on the framework AutoAttack. To avoid attack failure, we also used tools to verify the effectiveness of the attack. The empirical results from the experiments suggest that highly parameterized models are more robust than under-parameterized models. We discuss the background and the notations in Sect. 2, the methodology adopted to conduct experiments in Sect. 3, and experimental setup and results in Sect. 4. We conclude by discussing the major takeaways and future research directions in Sect. 5.
The code is available at https://github.com/pralab/overparam-adv
@@ 2.1 Over-parameterization and Double-descent curve
When designing a model architecture to solve an ML problem, from classical understanding, we choose a model that is rich enough to capture the underlying feature space without fitting too closely with the training data (or overfitting) and simple enough to avoid not being able to learn patterns in the data (or underfitting). The balance between underfitting and overfitting is widely known as bias-variance tradeoff plotted on a risk capacity curve as in Figure 1. From this understanding, we would want to find a ”sweet spot” where the model has enough capacity to perform well on test data while keeping the training error low. Having a large model with zero training error is conventionally avoided, as stated by Trevor Hastie in  [13], ”a model with zero training error is overfit to the training data and will typically generalize poorly” which is still a widely accepted notion. However, large NN are highly expressive and can efficiently optimize many parameters to fit the training data. Large number of parameters of the neural network, traditionally considered a case of overfitting, are better suited for generalization  [14]. Further, Belkin et al.’s  [15] who coined the term double-descent curve (left plot of Figure 1) suggested that when training a model on the increasing function capacity (referred to the size of network) until the point of zero training error/interpolation, typically have high risk and poor generalization error. Although, as explained in [16] it is not always a double descent curve, especially for deep neural networks, for which the bias usually dominates with respect to the variance and thus the risk is monotonically decreasing (right plot of Figure 1), unless there is noise in the training labels. Training beyond the interpolation point (in over-parameterized regime), it is observed that the variance of the model decreases, which leads to the overall decrease in the risk, giving improved generalization performance. Further work by Nakkiran et al. [17] showed that this generalization trend is not just a function of model capacity but also a function of number of epochs in training. In some literature, it is also argued that good generalization performance is due to the inductive bias of gradient-based algorithms, e.g. [18, 19].
@@ 2.2 Adversarial Robustness
$\textrm{s.t.}\quad\|\mathbf{x^{\prime}}-\mathbf{x}\|_{p}\leq\epsilon,$ (2)
@@ 2.3 Adversarial Robustness of Over-parameterized Network
While a considerable number of works [20, 21, 22, 23, 24] assess the robustness of adversarially trained networks, which is out of the scope of this article, the findings of various related works suggest that overparameterization would decrease robustness. In the seminal work by Fu et al. [25], the authors claim that decreasing the number of nodes reduces the sensitivity to small perturbations of the input samples. Sagawa et al. [26] suggested that the over-parameterization can cause the exacerbation of spurious correlation, i.e., the relation between the biased features of input like background, texture, or unimportant pixels with the labels. However, in [27], the authors showed that wider neural networks tend to have better robustness, assessed with the PGD attack. The results in this paper were also replicated in [28] and are supported by [29], where the authors claim that obtaining robustness may require exponentially more parameters than obtaining accuracy on the clean test data. The contradiction present in the literature has been highlighted in [30], which provides a theoretical analysis that only encompasses two-layer neural networks, and in [31], which, with an analysis limited to linear models, shows that over-parameterization may be either a source of additional robustness or brittleness. The analysis performed on deep neural networks by the works affirming that over-parameterization increases robustness [27, 28] is empirical and based on the PGD attack, and they have not employed any tool to check the reliability of the obtained results. Therefore, the attack might have failed [32, 33], providing unreliable results that have caused a false sense of robustness for the overparameterized networks. Because gradient-based attacks can fail in many cases [32, 33], e.g., if the employed hyperparameters are incorrect or the decision surface is not smooth enough, they can not achieve convergence, providing a suboptimal point that may not be able to evade the classifier. The goal of this work is to verify if this is the reason that led to the abovementioned contradiction between the state-of-the-art works. Analyzing the reliability of the empirical analysis to assess the robustness of overparametrized networks, we have demonstrated that it is reliable; thus, the higher robustness of overparametrized networks is not due to experimental problems.
@@ 3.1 Notations
We consider a data space $\mathcal{X}$, $\mathcal{Y}$] where $\mathcal{X}$ is a $d$-dimensional feature space of input samples and $\mathcal{Y}$ is a label space with c-classes such that $\mathcal{Y}\in[c]$. The training data is represented as $\mathcal{D}=(\mathbf{x}_{i},y_{i})_{i=1}^{n}$ where $n$ is the size of the training set. We use $L(\mathcal{D},\theta)$ to represent loss incurred by classifier $f\in\mathcal{H}:\mathcal{X}\rightarrow\mathcal{Y}$ where $\mathcal{H}$ is a class of function.Typically the loss is computed by averaging the sample-wise loss function $l(f(\mathbf{x}),y)$ over the training set, i.e., $\frac{1}{n}\sum_{i=1}^{n}l(f(\mathbf{x}_{i}),y_{i})$. We assume that the classifier $f$ is learned by minimizing the loss function and best function class $f^{*}$ is selected using empirical risk minimization:
@@ 3.2 Attack Approach
The attack approaches considered in our experiments have the following threat model: The attacker’s main goal is to cause integrity violation to evade detection without compromising normal operation using indiscriminate attack whose goal is to have the test-instance misclassified as any class except the true one. The attacker’s capability is limited to influencing data samples at test time to evade the trained model, making it an evasion attack. We consider both white-box setting, where the attacker has complete access to both model and data distribution, and the black-box setting, where the attacker has no access to either data or distribution. To evaluate the security of over-parameterized models, we use the widely-used white-box attack approach, which generates attacks on the known trained model parameters. PGD is a widely used evasion attack algorithm to evaluate the robustness of a model and is a good starting point for security evaluation. After evaluating on PGD, as explained in later sections, we perform experiments on sophisticated diversified ensemble attack: AutoAttack.
3.2.1 Projected Gradient Descent (PGD)
Proposed by Madry et al. [34], PGD is an iterative optimization-based attack widely used for the evaluation of the models’ robustness. This algorithm leverages Gradient Descent to compute a perturbation $\mathbf{\delta}$ on a sample $\mathbf{x}$ originally labeled as $y$, that maximizes the output of a loss function $\ell$ of choice, computed on the target model $f$. As illustrated in algorithm 1, at each iteration (line 1), PGD computed the gradient of the loss w.r.t the perturbation $\mathbf{\delta}$ (line 1), which is added to the previously-computed perturbation to maximize the error. The gradient is post-processed by considering only its sign, modulated by the step size $\alpha$. Then, the perturbation $\mathbf{\delta}$ is projected with $\prod$ inside the feasible domain $\Gamma$ (line 1). Such a domain defines both (i) the capability of the attacker, formalized as a bound $\epsilon$ on the p-norm of the perturbation itself resulting in $\parallel\mathbf{\delta}\parallel_{p}\leq\epsilon$ and (ii) the domain constraints needed to create valid and uncorrupted adversarial examples. At the end of the iterations, the perturbation that mostly maximizes the loss is added to the original sample to produce the adversarial example (line 1).
Input : $\mathbf{x},y$, the initial sample and its label; $f$, the target model; $\ell$ the loss function; $\prod_{\Gamma}$, the projection operator enforcing the constraints of the feasible domain $\Gamma$; $n$, the number of iterations; $\alpha$, the step size
Output : $\mathbf{x}^{\star}$, the solution found by the algorithm
1
[…]
3for $i\in[1,n]$ do
4 $\mathbf{\delta}_{i}\leftarrow\alpha\,sign(\nabla_{\mathbf{\delta}_{i}}\,\ell(f(\mathbf{x}+\mathbf{\delta}_{i}),y))$
5 $\mathbf{\delta}_{i}\leftarrow\prod_{\Gamma}(\mathbf{x},\mathbf{\delta}_{i})$
6 end for
[…]
Algorithm 1 Projected Gradient Descent.
@@ 3.2.2 AutoAttack
AutoAttack is an ensemble of different typologies of attacks [35]. It leverages three white-box and one black-box attack. The three white-box attacks include Automated PGD with Carlini-Wagner loss $APGD_{CW}$, Automated PGD with Difference of Logits Ratio loss $APGD_{DLR}$, targeted Fast Adaptive Boundary (FAB) attack, and Square Attacks as the black-box attack  [36]. The Automated PGD (APGD) is the standard PGD attack with a fixed number of iterations and adaptive step size. It uses Carlini-Wagner Loss and Difference of Logits Ratio loss instead of cross-entropy loss. $APGD_{CW}$ is an untargeted attack whereas $APGD_{DLR}$ is a targeted attack. Namely, it computes the adversarial example ad-hoc to have it misclassified as a specific class. If it does not succeed, it considers a different class as the target class. The third parameter-free white-box attack is the targeted version of the Fast Adaptive Boundary (FAB) attack  [37]. Finally, Square Attack is an untargeted black-box attack based on random search and is not affected by gradient masking. Gradient masking is a common adversarial defense [38] against gradient-based attack algorithms that makes the gradient usually computed by these attacks uninformative, which makes the attacker unable to compute an effective adversarial example. However, thanks to the transferability property of adversarial examples [39], these defense techniques can be easily broken by computing an attack against a model owned by the attacker (surrogate). These four attacks are packaged in a single module that serves as a diversified attack and helps evaluate the reliability of the results we obtained from our primary PGD experiments.
@@ 3.3 Evaluation Procedure
To evaluate the security of neural networks, we take into account different neural network architectures. These architectures are then scaled with different numbers of hidden units in their layers as done in [40]. We consider a model architecture, expanded to ten models ranging from under-parameterized to over-parameterized regime $\theta\in\{\theta_{1},\theta_{2},..,\theta_{m}\}$. The models are first trained on the training data. Then, for each model $\theta_{i}$, its base performance is evaluated using benign test data, i.e., where there is no perturbation $\epsilon_{0}$}. Consequently, we generate the adversarial examples ad-hoc against the considered model $\theta_{i}$ for varying degrees of attack budgets $\epsilon\in\{\epsilon_{1},\epsilon_{2}..,\epsilon_{k}\}$ and for each perturbation budget we compute the Error Rate (namely the percentage of misclassified samples) of $\theta_{i}$ on them. These error rates are then used to generate the SEC curve that shows how the error rate of the model $\theta_{i}$ increases for a growing perturbation budget of the attacker.